# MoD-SLAM: Monocular Dense Mapping for Unbounded 3D Scene Reconstruction

### 09 Feb 2024 v2

@@ 4.2 Results and Evaluation
Evaluation on Replica [28] Replica includes several geometric models reconstructed from real-world scenes. Table 2 shows the results of the reconstruction of systems on the Replica dataset. Compared to the systems mentioned in the table, our system exhibits superior tracking capabilities in both monocular and RGB-D modes. In comparison with NeRF-based SLAM systems[45, 42, 44], our approach achieves enhanced reconstruction fidelity. Compared to 3D Gaussian-based SLAM systems [11], our system falls in terms of scene reconstruction accuracy. However, our system outperforms in tracking performance and is more effective in recovering the scale of the scene. Fig. 4 and Fig. 6(a) present a comparative analysis of scene reconstructions by our system against NICE-SLAM [45] and GO-SLAM [42]. Our method demonstrates superior performance that shows smoother scene reconstruction and the ability to restore the texture and contours of geometric objects more accurately.
Evaluation on ScanNet [7] We evaluated our system across seven distinct scenes in ScanNet dataset [7]. Table 2 presents the tracking results. Our system outperforms others in tracking accuracy, particularly in the large indoor scenes of ScanNet. As illustrated in Fig. 5, our system outmatches GO-SLAM [42] by rendering the scene with enhanced detail fidelity, more accurately approximating the real-world geometry and ensuring smoother scene transitions. Furthermore, our framework possesses a pronounced proficiency in addressing scene sparsity through effective void interpolation.
[…]
1 0 0 0 0 31.33 15.08
1 1 0 0 0 85.19 26.76
1 1 1 0 0 85.42 26.88
1 1 1 1 0 88.03 27.59
1 1 1 1 1 88.73 27.84
ATE[cm]$\downarrow$ F-$\uparrow$ None 0.46 82.93 DepE. 0.39 86.66 DepE./DepD. 0.35 88.73 RGB-D 0.33 89.96
[…]
4.3 Ablation Study
@@
Unbounded Scene We also conduct tests to assess the impact of reparameterization Equ. 6 and Gaussian encoding Equ. 5 on reconstructing scenes with no defined boundaries. For comparison, we used NEWTON [14] as a baseline to evaluate our system’s performance. The quantitative data in Table 5 indicates that our reparameterization and Gaussian encoding significantly improve the accuracy in both scaling and geometric reconstruction of unbounded scenes. These approaches notably enhance our system’s overall performance. Our data were acquired using the RGB-D mode of our system within the ScanNet 0207 scene. NEWTON’s data are from the original paper [14]. The results on the ScanNet [7] indicate that our model shows greater robustness and more precise tracking abilities compared to NEWTON.
Figure 6: (a) shows the reconstruction result on Replica. (b) shows the run time and GPU memory used in SLAM systems.
4.4 Runtime Analysis
We test the tracking and the mapping time of NICER-SLAM [44], GO-SLAM [42] and our system using the same number of sampling points. Fig. 6(b) shows that our system is faster than GO-SLAM and NICER-SLAM in tracking process. However, because of the depth estimation module, our system is slightly slower than GO-SLAM in mapping process, but higher than NICER-SLAM. In addition, Fig. 6(b) also demonstrates that our model size is smaller and uses less memory. GO-SLAM and our system are benchmarked with an NVIDIA RTX 4090 GPU while NICER-SLAM is benchmarked with an A100.
5 Conclusion
We presented MoD-SLAM, a monocular implicit neural-based dense mapping method that improves both the speed and accuracy of NeRF-based SLAM systems. Our experiments demonstrate that by integrating a monocular depth estimation module, our system not only achieves a more accurate scene-scale reconstruction but also excels in reconstructing the details and textures of geometric objects compared to previous SLAM systems. Additionally, we show the superior performance of our system in reconstructing unbounded scenes after employing Gaussian encoding and reparametrization techniques.

### 21 Feb 2024 v3

@@ 4.2 Results and Evaluation
Evaluation on Replica [28] Table 1 shows the results of the reconstruction of systems on the Replica dataset. Compared to the systems mentioned in the table, our system exhibits superior tracking capabilities in both monocular and RGB-D modes. In comparison with NeRF-based SLAM [46, 43, 45], our approach achieves enhanced reconstruction fidelity. Compared to 3D Gaussian-based SLAM systems [11], our system falls in terms of scene reconstruction accuracy. However, our system outperforms in tracking performance and is more effective in recovering the scale of the scene. Fig. 4 presents a comparative analysis of scene reconstructions by our system against NICE-SLAM [46] and GO-SLAM [43]. Our method demonstrates superior performance that shows smoother scene reconstruction.
Evaluation on ScanNet [7] We evaluated our system across seven distinct scenes in ScanNet dataset [7]. Table 2 presents the tracking results. Our system outperforms others in tracking accuracy, particularly in the large indoor scenes of ScanNet. As illustrated in Fig. 6 and Fig. 5, our system outmatches GO-SLAM [43] by rendering the scene with enhanced detail fidelity, more accurately approximating the real-world geometry and ensuring smoother scene transitions. Furthermore, our framework possesses a pronounced proficiency in addressing scene sparsity through effective void interpolation.
5 Conclusion
We presented MoD-SLAM, a monocular implicit neural-based dense mapping method that demonstrates state-of-the-art accuracy in monocular reconstruction. Our experiments show that by integrating a monocular depth estimation module, our system not only achieves a more accurate scene-scale reconstruction but also excels in reconstructing the details and textures of geometric objects compared to previous SLAM systems. Additionally, we show the superior performance of our system in reconstructing unbounded scenes after employing Gaussian encoding and reparametrization techniques.
Appendix 0.A Appendix
0.A.1 Ablation Study
@@
Unbounded Scene We also conduct tests to assess the impact of reparameterization Equ. 6 and Gaussian encoding Equ. 5 on reconstructing scenes with no defined boundaries. For comparison, we used NEWTON [14] as a baseline to evaluate our system’s performance. The quantitative data in Table 5 indicates that our reparameterization and Gaussian encoding significantly improve the accuracy in both scaling and geometric reconstruction of unbounded scenes. These approaches notably enhance our system’s overall performance. Our data were acquired using the RGB-D mode of our system within the ScanNet 0207 scene. NEWTON’s data are from the original paper [14]. The results on the ScanNet [7] indicate that our model shows greater robustness and more precise mapping abilities compared to NEWTON.
Figure 7: (a) shows the reconstruction result on Replica [28]. (b) shows the run time and GPU memory used in SLAM systems.
[…]
1 0 0 0 0 34.33 15.78
1 1 0 0 0 85.79 27.76
1 1 1 0 0 85.98 27.89
1 1 1 1 0 89.01 28.13
1 1 1 1 1 89.73 28.37
ATE[cm]$\downarrow$ F-$\uparrow$
None 0.46 83.79
DepE. 0.39 87.96
DepE./DepD. 0.35 89.43
RGB-D 0.33 90.96
[…]
Con. 23.85 0.749
Con-Gau. 25.02 0.787
None - -
NEWTON[14] 23.15 0.702
0.A.2 Runtime Analysis
We test the tracking and the mapping time of NICER-SLAM [45], GO-SLAM [43] and our system using the same number of sampling points. Fig. 7(b) shows that our system is faster than GO-SLAM and NICER-SLAM in the tracking process. However, because of the depth estimation module, our system is slightly slower than GO-SLAM in the mapping process, but higher than NICER-SLAM. In addition, Fig. 7(b) also demonstrates that our model size is smaller and uses less memory. GO-SLAM and our system are benchmarked with an NVIDIA RTX 4090 GPU while NICER-SLAM is benchmarked with an A100.
Scene fr1-desk fr2-xyz fr3-office
GO-SLAM[43] 1.500 0.600 1.000
RGB-D Co-SLAM[38] 2.400 1.700 2.400
Ours 1.382 0.641 0.939
DROID-SLAM[34] 2.713 1.304 2.573
DeepV2D[33] 21.391 6.950 11.922
Mono ORB-SLAM3[4] 1.534 0.720 1.400
GO-SLAM[43] 2.381 1.093 1.773
Ours 1.481 0.686 1.101
Table 6: ATE RMSE [cm] on the TUM dataset [29]. The performance metrics reported by our system for each scene are derived from the average of three executions per scene.
0.A.3 Additional Experimental Results
TUM Dataset [29] We further test the tracking accuracy on the TUM dataset [29]. As shown in Table 6, Our method not only achieve the most accurate tracking performance in RGBD mode, but also outperforms classic SLAM systems as well as GO-SLAM [43] in monocular mode.
Replica Dataset [29] Here we show more results on Replica [29]. Fig. 8, Fig. 9 and Fig. 7(a) demonstrate the superior scene reconstruction capability of our proposed method in monocular and RGB-D mode, and also exhibiting enhanced smoothness in comparison to GO-SLAM [43].
Figure 8: Reconstruction results on Replica [28]. Qualitative comparison on Replica office4 and room2.
Figure 9: Reconstruction results on Replica [28]. Qualitative comparison on Replica office1 and office3.